delilahlouvre:

Hello, please help me solve this problem: The weight w (in pounds) that can be supported by a shelf made from half-inch Douglas fir plywood can be modeled by w=(82.9/d)3 where d is the distance (in inches) between the supports for the shelf. Find the inverse of the function. Then find the distance between the supports of a shelf that can hold set of encyclopedia weighing 66 pounds.

delilahlouvre:

@MrMudd183

delilahlouvre:

Anyone hellllppp :(( this is due tomorrow and I still cant figure it out.

delilahlouvre:

@supie

supie:

Is "w=(82.9/d)3 " supposed to say w=(82.9/d)^3 @delilahlouvre?

delilahlouvre:

delilahlouvre:

it's multiplied by 3 and not to ^3

supie:

So are we solving for d or w?

delilahlouvre:

find the inverse function of the model w= (82.9/d)3 first then after that solve for the distance between the supports of the shelf that can hold set of the encyclopedia weighing 66 pounds.

jhonyy9:

so find the inverse of the function mean \[w^{-1}(d) = ?\] do you know the way to calcule it ?

Mercury:

old question but will respond so this can be closed to find the inverse, swap your independent and dependent variable (w and d) and solve for w again. the result will be your inverse.